# Optimal and Fair Encouragement Policy Evaluation and Learning

### v4

@@ D.3.1 Preliminaries: results from other works used without proof
Definition 1 (Local Rademacher Complexity).
@@
Lemma 2 (Lemma 5 of [Chernozhukov et al., 2019]/Lemma 4, Foster and Syrgkanis [2019]).
@@
Lemma 3 (Concentration of conditional moments ([Agarwal et al., 2018, Woodworth et al., 2017]).
For any $j\in\mathcal{J}$, with probability at least $1-\delta$, for all $Q$,
$\left|\widehat{h}_{j}(Q)-h_{j}(Q)\right|\leq 2\mathcal{R}_{n_{j}}(\mathcal{Q})+\frac{2}{\sqrt{n_{j}}}+\sqrt{\frac{\ln(2/\delta)}{2n_{j}}}$
@@
$\left|\widehat{h}_{j}(Q)-h_{j}(Q)\right|\leq 2\mathcal{R}_{np_{j}^{\star}/2}(\mathcal{Q})+2\sqrt{\frac{2}{np_{j}^{\star}}}+\sqrt{\frac{\ln(4/\delta)}{np_{j}^{\star}}}$
Lemma 4 (Orthogonality (analogous to [Chernozhukov et al., 2019] (Lemma 8), others)).
@@
$V(Q_{0})-V(\hat{Q})\leq O(R_{n,\delta}+\chi^{2}_{n})$
D.4 Adapted lemmas
@@
Lemma 5 (Feasible vs. oracle nuisances in low-variance regret slices (Chernozhukov et al. [2019], Lemma 9) ).
Suppose that the mean squared error of the nuisance estimates is upper bounded w.p. $1-\delta$ by $\chi_{n,\delta}^{2}$ and suppose $\chi_{n,\delta}^{2}\leq\epsilon_{n}$. Then:
$V_{2}^{0}=\sup_{Q,Q^{\prime}\in\mathcal{Q}_{*}\left(\epsilon_{n}+2\chi_{n,\delta}^{2}\right)}\operatorname{Var}\left(v_{DR}^{0}(x;Q)-v_{DR}^{0}\left(x;Q^{\prime}\right)\right)$
@@
D.5 Proof of Theorem˜4
Proof of Theorem˜4.
We first study the meta-algorithm with “oracle" nuisance functions $\eta=\eta_{0}$. For brevity below we notationally suppress the dependence of $v$ on observation $O$.
Define
$\Pi_{2}\left(\epsilon_{n}\right)$ $=\left\{\pi\in{\Pi}:\mathbb{E}_{n_{1}}[v(\pi;\eta_{0})-v(\hat{Q}_{1};\eta_{0})]\leq\epsilon_{n},\mathbb{E}_{n_{1}}\left[g_{j}(O;\pi,\eta_{0})-g_{j}\left(O;\hat{Q}_{1},\eta_{0}\right)\mid\mathcal{E}_{j}\right]\leq\epsilon_{n},j\in\hat{\mathcal{I}}_{1}\right\}$
$\mathcal{Q}_{2}\left(\epsilon_{n}\right)$ $=\left\{Q\in\Delta\left({\Pi_{2}}(\epsilon_{n})\right)\right\}$
$\mathcal{Q}^{*}\left(\epsilon_{n}\right)$ $=\left\{Q\in\Delta({\Pi}):\mathbb{E}[(v(Q;\eta_{0})-v(Q^{*};\eta_{0})]\leq\epsilon_{n},\mathbb{E}[g_{j}(O;Q,\eta_{0})\mid\mathcal{E}_{j}]-\mathbb{E}[g_{j}(O;Q^{*},\eta_{0})\mid\mathcal{E}_{j}]\leq\epsilon_{n}\right\}$
@@ D.5 Proof of Theorem˜4
Invoking Theorem˜6 on the output of the first stage of the algorithm, yields that with probability $1-\frac{\delta}{6}$ over the randomness in $\mathcal{D}_{1},$ by choice of $\epsilon_{n}$ = $\bar{O}(n^{-\alpha}))$,
$V(\hat{Q}_{1})$ $\leq V(Q^{*})+\epsilon_{n}/2$
$\gamma_{k}(\hat{Q}_{1})$ $\leq d_{k}+\sum_{j\in\mathcal{J}}\left|M_{k,j}\right|\widetilde{O}\left((np_{j}^{*})^{-\alpha}\right)\leq d_{k}+\epsilon_{n}/2\quad\text{ for all }k$
Further, by Lemma˜2,
@@
$\sup_{Q\in\mathcal{Q}}\left|\mathbb{E}_{n_{1}}[(g(O;Q)-g(O;Q^{*}))]-\mathbb{E}[(g(O;Q)-g(O;Q^{*}))]\right|\leq\epsilon_{n}/2$
Therefore, with high probability on the good event, $Q^{*}\in\mathcal{Q}_{2}.$
Step 2: Again invoking Theorem˜6, this time on the output of the second stage of the algorithm with function space $\Pi_{2}$ (hence implicitly $\mathcal{Q}_{2}$), and conditioning on the “good event" that $Q^{*}\in\mathcal{Q}_{2}$, we obtain the bound that with probability $\geq 1-\delta/3$ over the randomness of the second sample $\mathcal{D}_{2},$
$V(\hat{Q}_{2})$ $\leq V(Q^{*})+\epsilon_{n}/2$
$\gamma_{k}(\hat{Q}_{2})$ $\leq\gamma_{k}({Q}^{*})+\epsilon_{n}/2$
@@
$\mathcal{Q}_{2}^{0}=\{Q\in\operatorname{conv}(\Pi)\colon\left|V(Q)-V(\hat{Q}_{1})\right|\leq\epsilon_{n}/2,\mathbb{E}[g_{j}(O;Q)-g_{j}(O;\hat{Q}_{1}))\mid\mathcal{E}_{j}]\leq\epsilon_{n},\forall j\}$
@@
$\sup_{Q\in\mathcal{Q}}\left|\mathbb{E}_{n_{1}}[g_{j}(O;Q)-g_{j}(O;\hat{Q}_{1})\mid\mathcal{E}_{j}]-\mathbb{E}[g_{j}(O;Q)-g_{j}(O;\hat{Q}_{1})\mid\mathcal{E}_{j}]\right|$
$\qquad\qquad\leq 2\sup_{Q\in\mathcal{Q}}\left|\mathbb{E}_{n_{1}}[g_{j}(O;Q)\mid\mathcal{E}_{j}]-\mathbb{E}[g_{j}(O;Q)\mid\mathcal{E}_{j}]\right|\leq\epsilon,\forall j$
@@
$\left|\mathbb{E}[g_{j}(O;Q)-g_{j}(O;Q^{*})\mid\mathcal{E}_{j}]\right|\leq\left|\mathbb{E}[g_{j}(O;Q)-g_{j}(O;\hat{Q}_{1})\mid\mathcal{E}_{j}]\right|+\left|\mathbb{E}[g_{j}(O;\hat{Q}_{1})-g_{j}(O;Q^{*})\mid\mathcal{E}_{j}]\right|\leq\epsilon_{n}$
Define the space of distributions over policies that achieve value and constraint regret in the population of at most $\epsilon_{n}:$
$\mathcal{Q}_{*}(\epsilon_{n})=\{Q\in\mathcal{Q}\colon V(Q)-V(Q^{*})\leq\epsilon_{n},\;\;\mathbb{E}[g_{j}(O;Q)-g_{j}(O;Q^{*})\mid\mathcal{E}_{j}]\leq\epsilon_{n},\forall j\},$
@@
$\mathcal{Q}_{2}^{0}(\epsilon_{n})\subseteq\mathcal{Q}_{*}(\epsilon_{n}).$ (10)
@@
$r_{2}^{2}$ $=\sup_{Q\in\mathcal{Q}_{2}}\mathbb{E}[(v(Q)-v(Q^{*}))^{2}]\leq\sup_{Q\in\mathcal{Q}^{*}(\epsilon_{n})}\mathbb{E}[(v(Q)-v(Q^{*}))^{2}]$
$=\sup_{Q\in\mathcal{Q}^{*}(\epsilon_{n})}\operatorname{Var}(v(Q)-v(Q^{*}))+\mathbb{E}[(v(Q)-v(Q^{*}))]^{2}$
$\leq\sup_{Q\in\mathcal{Q}^{*}(\epsilon_{n})}\operatorname{Var}(v(Q)-v(Q^{*}))+\epsilon_{n}^{2}$
@@
$r_{2}\leq\sqrt{\sup_{Q\in\mathcal{Q}_{*}\left(\epsilon_{n}\right)}\operatorname{Var}\left(v(Q)-v(Q^{*})\right)}+2\epsilon_{n}=\sqrt{V_{2}}+2\epsilon_{n}$
Combining this with the local Rademacher complexity bound, we obtain that:
$\mathbb{E}[v(\hat{Q}_{2})-v(Q^{*})]=O\left(\kappa\left(\sqrt{V_{2}}+2\epsilon_{n},\mathcal{Q}_{*}\left(\epsilon_{n}\right)\right)+\sqrt{\frac{V_{2}\log(3/\delta)}{n}}\right)$
These same arguments apply for the variance of the constraints
$V_{2}^{j}=\sup\;\{\operatorname{Var}\left(g_{j}(O;Q)-g_{j}\left(O;Q^{\prime}\right)\right)\colon{Q,Q^{\prime}\in\mathcal{Q}_{*}(\tilde{\epsilon}_{n})}\}$
D.6 Proofs of auxiliary/adapted lemmas
Proof of Lemma˜5.
The proof is analogous to that of [Chernozhukov et al., 2019, Lemma 9] except for the step of establishing that $\pi_{*}\in\mathcal{Q}^{0}_{\epsilon_{n}+O(\chi_{n,\delta}^{2})}$: in our case we must establish relationships between saddlepoints under estimated vs. true nuisances. We show an analogous version below.
@@ D.6 Proofs of auxiliary/adapted lemmas
$(Q^{*}_{0,0},\lambda^{*}_{0,0})$ $\in\arg\min_{Q}\max_{\lambda}\mathbb{E}[v_{DR}(Q;\eta_{0})]+\lambda^{\top}(\gamma_{DR}(Q;\eta_{0})-d)\coloneqq L(Q,\lambda;\eta_{0},\eta_{0})\coloneqq L(Q,\lambda),$
$(Q^{*}_{\eta,0},\lambda^{*}_{\eta,0})$ $\in\arg\min_{Q}\max_{\lambda}\mathbb{E}[v_{DR}(Q;\eta)]+\lambda^{\top}(\gamma_{DR}(Q;\eta_{0})-d),$
$(Q^{*},\lambda^{*})$ $\in\arg\min_{Q}\max_{\lambda}\mathbb{E}[v_{DR}(Q;\eta)]+\lambda^{\top}(\gamma_{DR}(Q;\eta)-d).$
@@
$\mathbb{E}[v_{DR}(Q^{*})]\leq L(Q^{*},\lambda^{*};\eta,\eta)+\nu$
$\leq L(Q^{*},\lambda^{*};\eta,\eta_{0})+\nu+\chi_{n,\delta}^{2}$
$\leq L(Q^{*},\lambda^{*};\eta,\eta_{0})+\nu+\chi_{n,\delta}^{2}$ by Lemma˜4
$\leq L(Q^{*},\lambda^{*}_{\eta,0};\eta,\eta_{0})+\nu+\chi_{n,\delta}^{2}$ by saddlepoint prop.
$\leq L(Q^{*}_{\eta,0},\lambda^{*}_{\eta,0};\eta,\eta_{0})+\left|L(Q^{*}_{\eta,0},\lambda^{*}_{\eta,0};\eta,\eta_{0})-L(Q^{*},\lambda^{*}_{\eta,0};\eta,\eta_{0})\right|+\nu+\chi_{n,\delta}^{2}$ triangle ineq.
$\leq L(Q^{*}_{\eta,0},\lambda^{*}_{\eta,0};\eta,\eta_{0})+\epsilon_{n}+\nu+\chi_{n,\delta}^{2}$ $\text{assuming }\epsilon_{n}\geq\chi_{n,\delta}^{2}$
$\leq\mathbb{E}[v_{DR}(Q^{*}_{\eta,0};\eta)]+\epsilon_{n}+2\nu+\chi_{n,\delta}^{2}$ apx. complementary slackness
$\leq\mathbb{E}[v_{DR}(Q^{*}_{0,0};\eta)]+\epsilon_{n}+2\nu+\chi_{n,\delta}^{2}$ suboptimality
@@
$\mathbb{E}[v_{DR}(Q^{*};\eta)]-\mathbb{E}[v_{DR}(Q^{*}_{0,0};\eta)]\leq\epsilon_{n}+2\nu+\chi_{n,\delta}^{2}.$
@@
Applying Lemma˜4 gives;
$V(Q^{*})-V(Q^{*}_{0,0})\leq\epsilon_{n}+2\nu+2\chi_{n,\delta}^{2}.$
Define policy classes with respect to small-population regret slices (with a nuisance-estimation enlarged radius):
$\mathcal{Q}^{0}(\epsilon)=\{Q\in\Delta(\Pi)\colon V(Q^{*}_{0})-V(Q)\leq\epsilon,\gamma(Q^{*}_{0})-\gamma(Q)\leq\epsilon\}$
@@
$V_{2}^{obj}\leq\sup_{Q\in\mathcal{Q}^{0}(\epsilon_{n})}\operatorname{Var}(v_{DR}(O;\pi)-v_{DR}(O;\pi^{*})),$
where we have shown that $\pi^{*}\in\mathcal{Q}^{0}(\epsilon+2\nu+2\chi_{n,\delta}^{2}).$
Following the rest of the argumentation in [Chernozhukov et al., 2019, Lemma 9] from here onwards gives the result, i.e. studying the case of estimated nuisances with our Lemma˜5 and Lemma˜4. ∎
Appendix E Additional case study
E.1 Text message reminders

### v5

@@ 12.4.1 Preliminaries: results from other works used without proof
Definition 12.13 (Local Rademacher Complexity)
@@
Lemma 12.14 (Lemma 5 of [Chernozhukov et al., 2019]/Lemma 4, Foster and Syrgkanis [2019])
@@
Lemma 12.15 (Concentration of empirical constraint moments ([Agarwal et al., 2018, Woodworth et al., 2017]))
For any $k\in[K]$, with probability at least $1-\delta$, for all $Q$,
$\left|\widehat{\Delta}_{k}(Q)-\Delta_{k}(Q)\right|\leq 2\mathcal{R}_{n_{j}}(\mathcal{Q})+\frac{2}{\sqrt{n_{j}}}+\sqrt{\frac{\ln(2/\delta)}{2n_{j}}}$
@@
$\left|\widehat{\Delta}_{k}(Q)-\Delta_{k}(Q)\right|\leq 2\mathcal{R}_{np_{j}^{\star}/2}(\mathcal{Q})+2\sqrt{\frac{2}{np_{j}^{\star}}}+\sqrt{\frac{\ln(4/\delta)}{np_{j}^{\star}}}$
Lemma 12.16 (Orthogonality (analogous to [Chernozhukov et al., 2019] (Lemma 8), others))
@@
$V(Q_{0})-V(\hat{Q})\leq O(R_{n,\delta}+\chi^{2}_{n,\delta})$
12.5 Adapted lemmas
@@
Lemma 12.17 (Feasible vs. oracle nuisances in low-variance regret slices (Chernozhukov et al. [2019], Lemma 9) )
Suppose that the mean squared error of the nuisance estimates is upper bounded w.p. $1-\delta$ by $\chi^{2}_{n,\delta}$ and suppose $\chi^{2}_{n,\delta}\leq\epsilon_{n}$. Then:
$V_{2}^{0}=\sup_{Q,Q^{\prime}\in\mathcal{Q}^{\star}\left(\epsilon_{n}+2\chi^{2}_{n,\delta};\hat{I}_{1}\right)}\operatorname{Var}\left(v_{DR}^{0}(x;Q)-v_{DR}^{0}\left(x;Q^{\prime}\right)\right)$
@@
12.6 Proof of Theorem 5.6
Proof 12.18
Proof of Theorem 5.6
We first study the meta-algorithm with “oracle” nuisance functions $\eta=\eta_{0}$. For brevity below we notationally suppress the dependence of $v$ on observation $O$.
Define the empirical second-stage feasible set:
$\mathcal{Q}_{2}(\epsilon_{n}):=\left\{Q\in\mathcal{P}(\Pi):\mathbb{E}_{n_{1}}\!\left[v_{\mathrm{DR}}(O;\hat{Q}_{1},\eta_{0})-v_{\mathrm{DR}}(O;Q,\eta_{0})\right]\leq\epsilon_{n},\quad\mathbb{E}_{n_{1}}\!\left[\delta_{k}(O;Q,\eta_{0})-\delta_{k}(O;\hat{Q}_{1},\eta_{0})\right]\leq\epsilon_{n},\ \forall k\in\hat{I}_{1},\quad\mathbb{E}_{n_{1}}[\delta_{k}(O;Q,\eta_{0})]\leq\hat{d}_{k},\ \forall k\in[K]\right\}.$
By construction, the second-stage optimization localizes only the nearly binding constraints indexed by $\hat{I}_{1}$, while retaining the slackened feasibility constraints for all $k\in[K]$.
@@
Invoking Theorem 12.12 on the output of the first stage of the algorithm, yields that with probability $1-\frac{\delta}{6}$ over the randomness in $\mathcal{D}_{1},$ by choice of $\epsilon_{n}$ = $\bar{O}(n^{-\alpha}))$,
$V(\hat{Q}_{1})$ $\geq V(Q^{*})-\epsilon_{n}/2$
$\Delta_{k}(\hat{Q}_{1})$ $\leq d_{k}+\sum_{j\in\mathcal{J}}\left|M_{k,j}\right|\widetilde{O}\left((np_{j}^{*})^{-\alpha}\right)\leq d_{k}+\epsilon_{n}/2\quad\text{ for all }k$
Further, by Lemma 12.14,
@@
$\sup_{Q\in\mathcal{Q}}\left|\mathbb{E}_{n_{1}}[(g(O;Q)-g(O;Q^{*}))]-\mathbb{E}[(\delta_{k}(O;Q)-\delta_{k}(O;Q^{*}))]\right|\leq\epsilon_{n}/2$
Therefore, with high probability on the good event, $Q^{\star}\in\mathcal{Q}_{2}(\epsilon_{n})$.
Step 2: Step 2: Again invoking Theorem 12.12 on the output of the second stage of the algorithm, restricted to the localized feasible class $\mathcal{Q}_{2}(\epsilon_{n})$, and conditioning on the good event that $Q^{\star}\in\mathcal{Q}_{2}(\epsilon_{n})$, we obtain that with probability at least $1-\delta/3$ over the randomness of the second sample $\mathcal{D}_{2}$,
$V(\hat{Q}_{2})$ $\geq V(Q^{*})-\epsilon_{n}/2$
$\Delta_{k}(\hat{Q}_{2})$ $\leq\Delta_{k}({Q}^{*})+\epsilon_{n}/2$
@@
$\mathcal{Q}_{2}^{0}=\{Q\in\operatorname{conv}(\Pi)\colon\left|V(Q)-V(\hat{Q}_{1})\right|\leq\epsilon_{n}/2,\mathbb{E}[\delta_{k}(O;Q)-\delta_{k}(O;\hat{Q}_{1})]\leq\epsilon_{n},\forall k\}$
@@
$\sup_{Q\in\mathcal{Q}}\left|\mathbb{E}_{n_{1}}[\delta_{k}(O;Q)-\delta_{k}(O;\hat{Q}_{1})]-\mathbb{E}[\delta_{k}(O;Q)-\delta_{k}(O;\hat{Q}_{1})]\right|$
$\qquad\qquad\leq 2\sup_{Q\in\mathcal{Q}}\left|\mathbb{E}_{n_{1}}[\delta_{k}(O;Q)]-\mathbb{E}[\delta_{k}(O;Q)]\right|\leq\epsilon,\forall j$
@@
$\left|\mathbb{E}[\delta_{k}(O;Q)-\delta_{k}(O;Q^{*})]\right|\leq\left|\mathbb{E}[\delta_{k}(O;Q)-\delta_{k}(O;\hat{Q}_{1})]\right|+\left|\mathbb{E}[\delta_{k}(O;\hat{Q}_{1})-\delta_{k}(O;Q^{*})]\right|\leq\epsilon_{n}$
Define the population localized near-optimal slice:
$\mathcal{Q}^{\star}(\epsilon_{n};\hat{I}_{1}):=\left\{Q\in\mathcal{P}(\Pi):V(Q^{\star})-V(Q)\leq\epsilon_{n},\quad\Delta_{k}(Q)\leq d_{k}+\epsilon_{n},\ \forall k\in[K],\quad\Delta_{k}(Q)-\Delta_{k}(Q^{\star})\leq\epsilon_{n},\ \forall k\in\hat{I}_{1}\right\}.$
On the same high-probability event, the population counterpart of the empirical second-stage feasible set satisfies
$\mathcal{Q}_{2}(\epsilon_{n})\subseteq\mathcal{Q}^{\star}(C\epsilon_{n};\hat{I}_{1})$
for a universal constant $C<\infty$. The value part follows from
$V(Q^{\star})-V(Q)=\{V(Q^{\star})-V(\hat{Q}_{1})\}+\{V(\hat{Q}_{1})-V(Q)\}\leq C\epsilon_{n},$
while the full feasibility constraints follow from the slackened empirical constraints and uniform concentration:
$\Delta_{k}(Q)\leq d_{k}+C\epsilon_{n},\qquad\forall k\in[K].$
For the localized comparison constraints, the definition of $\hat{I}_{1}$, the first-stage near-binding property, and concentration imply
$\Delta_{k}(Q)-\Delta_{k}(Q^{\star})\leq C\epsilon_{n},\qquad\forall k\in\hat{I}_{1}.$
@@
$\mathcal{Q}_{2}(\epsilon_{n})\subseteq\mathcal{Q}^{\star}(C\epsilon_{n};\hat{I}_{1}).$ (13)
@@
$r_{2}^{2}=\sup_{Q\in\mathcal{Q}_{2}}\mathbb{E}[(v(Q)-v(Q^{\star}))^{2}]\leq\sup_{Q\in\mathcal{Q}^{\star}(C\epsilon_{n};\hat{I}_{1})}\mathbb{E}[(v(Q)-v(Q^{\star}))^{2}]$
$\sup_{Q\in\mathcal{Q}^{\star}(C\epsilon_{n};\hat{I}_{1})}\operatorname{Var}(v(Q)-v(Q^{\star}))+\mathbb{E}[(v(Q)-v(Q^{\star}))]^{2}$
$\leq\sup_{Q\in\mathcal{Q}^{\star}(C\epsilon_{n};\hat{I}_{1})}\operatorname{Var}(v(Q)-v(Q^{\star}))+\epsilon_{n}^{2}.$
@@
$r_{2}\leq\sqrt{\sup_{Q\in\mathcal{Q}^{\star}(C\epsilon_{n};\hat{I}_{1})}\operatorname{Var}\left(v(Q)-v(Q^{\star})\right)}+2\epsilon_{n}=\sqrt{V_{2}}+2\epsilon_{n}$
Therefore, applying the local Rademacher bound on the localized class $\mathcal{Q}^{\star}(C\epsilon_{n};\hat{I}_{1})$ gives
$V(Q^{\star})-V(\hat{Q}_{2})=O\!\left(\kappa(\bar{\sigma}_{D_{2}},\operatorname{conv}(\mathcal{F}_{\Pi}))+\bar{\sigma}_{D_{2}}n_{2}^{-1/2}\sqrt{\log(3/\delta)}+\chi^{2}_{n,\delta}\right).$
Applying the same argument coordinatewise to the constraint scores gives, for each $k\in[K]$,
$\bigl(\Delta_{k}(\hat{Q}_{2})-d_{k}\bigr)_{+}=O\!\left(\kappa(\bar{\sigma}_{k,D_{2}},\operatorname{conv}(\mathcal{F}^{\Delta}_{k}))+\bar{\sigma}_{k,D_{2}}n_{2}^{-1/2}\sqrt{\log(3/\delta)}+\chi^{2}_{n,\delta}\right).$
12.7 Proofs of auxiliary/adapted lemmas
Proof 12.19
Proof of Lemma 12.17 The proof is analogous to that of [Chernozhukov et al., 2019, Lemma 9] except for the step of establishing that $\pi_{*}\in\mathcal{Q}^{0}_{\epsilon_{n}+O(\chi_{n,\delta}^{2})}$: in our case we must establish relationships between saddlepoints under estimated vs. true nuisances. We show an analogous version below.
For this proof only, switch to the equivalent cost formulation
$\tilde{v}_{\mathrm{DR}}(O;Q,\eta):=-v_{\mathrm{DR}}(O;Q,\eta),\qquad\tilde{V}(Q):=-V(Q).$
Thus the minimization saddle-point problem below is the costized version of the main text’s value-maximization problem.
@@ 12.7 Proofs of auxiliary/adapted lemmas
$(Q^{*}_{0,0},\lambda^{*}_{0,0})$ $\in\arg\min_{Q}\max_{\lambda}\mathbb{E}[\tilde{v}_{\mathrm{DR}}(Q;\eta_{0})]+\lambda^{\top}(\Delta(Q;\eta_{0})-d)=:\mathcal{L}(Q,\lambda;\eta_{0},\eta_{0})=:\mathcal{L}(Q,\lambda),$
$(Q^{*}_{\eta,0},\lambda^{*}_{\eta,0})$ $\in\arg\min_{Q}\max_{\lambda}\mathbb{E}[\tilde{v}_{\mathrm{DR}}(Q;\eta)]+\lambda^{\top}(\Delta(Q;\eta_{0})-d),$
$(Q^{*},\lambda^{*})$ $\in\arg\min_{Q}\max_{\lambda}\mathbb{E}[\tilde{v}_{\mathrm{DR}}(Q;\eta)]+\lambda^{\top}(\Delta(Q;\eta)-d).$
@@
$\mathbb{E}[\tilde{v}_{\mathrm{DR}}(Q^{*};\eta)]\leq\mathcal{L}(Q^{*},\lambda^{*};\eta,\eta)+\nu$
$\leq\mathcal{L}(Q^{*},\lambda^{*};\eta,\eta_{0})+\nu+\chi_{n,\delta}^{2}$
$\leq\mathcal{L}(Q^{*},\lambda^{*};\eta,\eta_{0})+\nu+\chi_{n,\delta}^{2}$ by Lemma 12.16
$\leq\mathcal{L}(Q^{*},\lambda^{*}_{\eta,0};\eta,\eta_{0})+\nu+\chi_{n,\delta}^{2}$ by saddlepoint prop.
$\leq\mathcal{L}(Q^{*}_{\eta,0},\lambda^{*}_{\eta,0};\eta,\eta_{0})+\left|\mathcal{L}(Q^{*}_{\eta,0},\lambda^{*}_{\eta,0};\eta,\eta_{0})-\mathcal{L}(Q^{*},\lambda^{*}_{\eta,0};\eta,\eta_{0})\right|+\nu+\chi_{n,\delta}^{2}$ triangle ineq.
$\leq\mathcal{L}(Q^{*}_{\eta,0},\lambda^{*}_{\eta,0};\eta,\eta_{0})+\epsilon_{n}+\nu+\chi_{n,\delta}^{2}$ $\text{assuming }\epsilon_{n}\geq\chi_{n,\delta}^{2}$
$\leq\mathbb{E}[\tilde{v}_{\mathrm{DR}}(Q^{*}_{\eta,0};\eta)]+\epsilon_{n}+2\nu+\chi_{n,\delta}^{2}$ apx. complementary slackness
$\leq\mathbb{E}[\tilde{v}_{\mathrm{DR}}(Q^{*}_{0,0};\eta)]+\epsilon_{n}+2\nu+\chi_{n,\delta}^{2}$ suboptimality
@@
$\tilde{V}(Q^{*})-\tilde{V}(Q^{*}_{0,0})\leq\epsilon_{n}+2\nu+\chi_{n,\delta}^{2}.$
Equivalently, since $\tilde{V}=-V$,
$V(Q^{*}_{0,0})-V(Q^{*})\leq\epsilon_{n}+2\nu+\chi_{n,\delta}^{2}.$
@@
Applying Lemma 12.16 gives:
$V(Q^{*}_{0,0})-V(Q^{*})\leq\epsilon_{n}+2\nu+2\chi_{n,\delta}^{2}.$
Define policy classes with respect to localized small-population regret slices (with a nuisance-estimation enlarged radius):
$\mathcal{Q}^{0}(\epsilon;\hat{I}_{1})=\left\{Q\in\mathcal{P}(\Pi):V(Q^{\star}_{0,0})-V(Q)\leq\epsilon,\quad\Delta_{k}(Q)\leq d_{k}+\epsilon,\ \forall k\in[K],\quad\Delta_{k}(Q)-\Delta_{k}(Q^{\star}_{0,0})\leq\epsilon,\ \forall k\in\hat{I}_{1}\right\}.$
@@
$V_{2}^{obj}\leq\sup_{Q\in\mathcal{Q}^{0}(\epsilon_{n};\hat{I}_{1})}\operatorname{Var}(v_{DR}(O;\pi)-v_{DR}(O;\pi^{*})),$
where we have shown that $\pi^{*}\in\mathcal{Q}^{0}(\epsilon+2\nu+2\chi_{n,\delta}^{2};\hat{I}_{1}).$
Following the rest of the argumentation in [Chernozhukov et al., 2019, Lemma 9] from here onwards gives the result, i.e. studying the case of estimated nuisances with our Lemma 12.17 and Lemma 12.16.
13 Additional case studies and additional details on experiments
13.1 SFHSA - SNAP case study
13.1.1 Additional descriptive statistics